# An overlooked source of uncertainty in the mass of the Milky Way

### v1

@@ 1 Correlations across radii in rotation curve measurements
Rotation curves are a widely-used dynamical tracer of the mass content of and distribution within galaxies including the Milky Way (e.g. Carignan et al., 2006; de Blok et al., 2008; Kuzio de Naray et al., 2008; Swaters et al., 2009; McMillan, 2011; Bovy et al., 2012; Sofue, 2012; Adams et al., 2014; Pato et al., 2015; Lelli et al., 2016; Eilers et al., 2019; Cautun et al., 2020; Ou et al., 2024, amongst many others). The atomic measurements comprising a rotation curve – the orbital speed at fixed radius – are usually taken to be independent from and uncorrelated with their counterparts at other radii (including in all the references above; see also Põder et al., 2023, for some discussion of concerns around such correlations). However, it is easy to see that correlations across radii must exist. One unambiguous source is the connection between gravitationally-driven kinematics and integrals of the mass distribution – for example, the mass within some central aperture contributes to the kinematics at all larger radii, introducing a correlation. Correlations may also arise due to instrumental effects such as beam smearing (Swaters et al., 2009, and references therein), or physical effects such as a spiral arm coherently perturbing the kinematics over a range in radius. The presence of correlations can further be inferred by noticing that the scatter implied by the statistical uncertainties on rotation curve measurements often exceeds the point-to-point scatter measured (but misestimates of the uncertainties could also contribute).
@@ 3 Impact of correlation modeling on inferred Milky Way mass profiles
Figure 1: Illustrative mass models for the Milky Way, with and without modeling radial correlations. The measurements (points with $1\sigma$ error bars) are as presented in Ou et al. (2024). In all cases the stellar disc (dashed yellow) and bulge (dotted yellow; combined stellar components shown with solid yellow), and ‘gas’ (green; including H i gas, H2 gas, warm dust and cold dust) are kept fixed to the model proposed by Ou et al. (2024); see Sec. 2 for details. In the left panels the halo component (solid purple) is an NFW model, while in the right panels it is a pseudo-isothermal sphere. In the upper panels no correlation (‘no GP’) between the observed rotation speeds is assumed, while in the lower panels the covariance matrix for the observations is estimated using Gaussian process (‘GP’) regression as described by Posti (2022). Shaded bands mark the regions enclosing 95 per cent of model curves at each radius for the halo component and model total.
@@ Acknowledgements
We thank Lorenzo Posti for making a well-documented software implementation of the methods presented in Posti (2022) available. KAO acknowledges support support by the Royal Society through Dorothy Hodgkin Fellowship DHF/R1/231105, by STFC through grant ST/T000244/1, and by the European Research Council (ERC) through an Advanced Investigator grant to C.S. Frenk, DMIDAS (GA 786910). AHR is supported by a Research Fellowship from the Royal Commission for the Exhibition of 1851. This research has made use of NASA’s Astrophysics Data System.

### v2

@@ 1 Correlations across radii in rotation curve measurements
Rotation curves are a widely-used dynamical tracer of the mass content of and distribution within galaxies including the Milky Way (e.g. Carignan et al., 2006; de Blok et al., 2008; Kuzio de Naray et al., 2008; Swaters et al., 2009; McMillan, 2011; Bovy et al., 2012; Sofue, 2012; Adams et al., 2014; Pato et al., 2015; Lelli et al., 2016; Eilers et al., 2019; Cautun et al., 2020; Ou et al., 2024, amongst many others). The atomic measurements comprising a rotation curve – the orbital speed at fixed radius – are usually taken to be independent from and uncorrelated with their counterparts at other radii (including in all the references above; see also Põder et al., 2023, for some discussion of concerns around such correlations). However, it is easy to see that correlations across radii must exist. One unambiguous source is the connection between gravitationally-driven kinematics and integrals of the mass distribution – for example, the mass within some central aperture contributes to the kinematics at all larger radii, introducing a correlation. Correlations may also arise due to instrumental effects such as beam smearing (Swaters et al., 2009, and references therein), modelling effects such as a geometrically thick disc being imperfectly separated into rings, or physical effects such as a spiral arm coherently perturbing the kinematics over a range in radius. The presence of correlations can further be inferred by noticing that the scatter implied by the statistical uncertainties on rotation curve measurements often exceeds the point-to-point scatter measured (but misestimates of the uncertainties could also contribute, including incorrectly assuming that uncertainties are Gaussian distributed).
@@ 3 Impact of correlation modeling on inferred Milky Way mass profiles
Figure 1: Illustrative mass models for the Milky Way, with and without modeling radial correlations. The measurements (points with $1\sigma$ error bars) are as presented in Ou et al. (2024). In all cases the stellar disc (dashed yellow) and bulge (dotted yellow; combined stellar components shown with solid yellow), and ‘gas’ (green; including H i gas, H2 gas, warm dust and cold dust) are kept fixed to the model proposed by Ou et al. (2024); see Sec. 2 for details. In the left panels the halo component (solid purple) is an NFW model, while in the right panels it is a pseudo-isothermal sphere. In the upper panels no correlation (‘no GP’) between the observed rotation speeds is assumed, while in the lower panels the covariance matrix for the observations is estimated using Gaussian process (‘GP’) regression as described by Posti (2022). Shaded bands mark the regions enclosing 95 per cent of model curves at each radius for the halo component and model total (in the upper panels these bands are very thin).
@@ 4.2 Importance of statistical uncertainty in the uncertainty budget
4.3 Possible pitfalls in interpretation
The approach of modelling correlations in rotation curve measurements across radii by adding degrees of freedom ($a_{k}$ and $s_{k}$) to the model and marginalizing over them requires care in interpretation. We have experimented with applying the same methodology to external galaxies and have encountered cases where the proposed mass model is unable to describe the data (an NFW dark halo being fit to a linearly rising rotation curve, for example). In such cases the parameter search responds by moving to very large values of $a_{k}$ (e.g. $>1000\,\mathrm{km}\,\mathrm{s}^{-1}$) and $s_{k}$ (e.g. $>100\,\mathrm{kpc}$), bounded above only by the boundary imposed on the (flat) prior. Such cases are clear model failures, but serve to illustrate that the parameters $a_{k}$ and $s_{k}$ will respond to essentially any discrepancy between the model and data by proposing a stronger correlation. In our exploration of the Milky Way, it is clear that something is missing from the models because $\chi^{2}_{\nu}\gg 1$ (see Figs. 4 & 5). Allowing for correlations in the measurements provides a plausible extension to the models: the goodness of fit improves and the additional parameters converge to plausible values corresponding to a fraction of the rotation speed and linear size of the disc. The question of whether this provides a more compelling explanation than other ways of accounting for the discrepancy between the data and models when correlations are neglected remains.
It is very challenging to account for all possible sources of correlation. Attempting to write down the full covariance matrix for a rotation curve measurement is currently infeasible – indeed this provides the motivation for the approach of Posti (2022). However, taking some burden of capturing correlations off of the model by describing them in the data and its associated uncertainties would undoubtedly help to mitigate the possible pitfalls described above.
@@ 5 Summary
  • it also results in larger uncertainty estimates for mass model parameters that we feel are more representative of the constraining power of the data;
@@ Acknowledgements
We thank Lorenzo Posti for making a well-documented software implementation of the methods presented in Posti (2022) available. We thank the anonymous referee for a constructive report. KAO acknowledges support support by the Royal Society through Dorothy Hodgkin Fellowship DHF/R1/231105, by STFC through grant ST/T000244/1, and by the European Research Council (ERC) through an Advanced Investigator grant to C.S. Frenk, DMIDAS (GA 786910). AHR is supported by a Research Fellowship from the Royal Commission for the Exhibition of 1851. This research has made use of NASA’s Astrophysics Data System.